A capacitor is charged to a potential of 12.0 V and then connected to a voltmeter having an

internal resistance of 3.40 MΩ. After a time of 4.00 s the voltmeter reads 3.0 V. What are

(a) the capacitance and (b) the time constant of the circuit?


1
Expert's answer
2022-03-30T13:51:07-0400

We know that

"Q=q_0e^\frac{-t}{RC}"

Q=CV

"V=V_0e^\frac{-t}{RC}"

"V=12\times e^\frac{-4}{3.40\times10^6\times C}"

(2)

Time constant

"\tau =RC"

"\tau=3.40\times10^6\times C"